Please help me to solve each equation for y:
8
Thank you very much

Answer by malakumar_kos@yahoo.com(315)


) y-1/x+3=-3/4 (solve)

y = -3/4+1/x-3 = -3x+4-12x/4x = -15x+4/4x
sol'n for y in terms of x = -15x+4/4x.